# EasyRec: Simple yet Effective Language Models for Recommendation

### v3

@@ Abstract.
Deep neural networks have become a powerful technique for learning representations from user-item interaction data in collaborative filtering (CF) for recommender systems. However, many existing methods heavily rely on unique user and item IDs, which limits their ability to perform well in practical zero-shot learning scenarios where sufficient training data may be unavailable. Inspired by the success of language models (LMs) and their strong generalization capabilities, a crucial question arises: How can we harness the potential of language models to empower recommender systems and elevate its generalization capabilities to new heights? In this study, we propose EasyRec - an effective and easy-to-use approach that seamlessly integrates text-based semantic understanding with collaborative signals. EasyRec employs a text-behavior alignment framework, which combines contrastive learning with collaborative language model tuning, to ensure a strong alignment between the text-enhanced semantic space and the collaborative behavior information. Extensive empirical evaluations across diverse real-world datasets demonstrate the superior performance of EasyRec compared to state-of-the-art alternative models, particularly in the challenging text-based zero-shot recommendation scenarios. Furthermore, the study highlights the potential of seamlessly integrating EasyRec as a plug-and-play component into text-enhanced collaborative filtering frameworks, thereby empowering existing recommender systems to elevate their recommendation performance and adapt to the evolving user preferences in dynamic environments. For better result reproducibility of our EasyRec framework, the model implementation details, source code, and datasets are available at the link: https://github.com/HKUDS/EasyRec.
copyright: none
1. Introduction
@@
Deep learning has established itself as a highly promising and powerful solution for capturing user preferences in the context of online recommender systems (Yang et al., 2022; Zhang et al., 2022). This approach harnesses the power of deep neural networks to learn rich and meaningful user and item representations by analyzing the complex patterns of user-item interactions. This, in turn, enables recommendation algorithms to accurately infer user preferences and provide highly relevant and personalized recommendations (Xu et al., 2023; Sun et al., 2021).
In recent years, there has been a notable surge of advancements in enhancing recommender systems through the incorporation of neural network-powered collaborative filtering frameworks, particularly in the domain of graph neural networks (GNNs) (Wang et al., 2019; He et al., 2020; Xia et al., 2022). By effectively leveraging the inherent graph structure present in the data, GNNs have demonstrated exceptional capabilities in capturing high-order relationships and complex dependencies among users and items. Notable examples of such GNN-based approaches include NGCF (Wang et al., 2019) and LightGCN (He et al., 2020). These methods have showcased impressive performance in recommendation tasks by virtue of their ability to model the collaborative signals present in the data through recursive message passing mechanisms.
The issue of data scarcity in recommender systems poses a significant challenge for existing deep collaborative filtering models, hindering their ability to learn precise user/item representations, particularly when dealing with sparse interaction data (Lin et al., 2021; Wei et al., 2021; Ren et al., 2024b; Hao et al., 2021). This challenge primarily arises from the limited availability of training labels, which in turn impedes the models’ capacity to capture the intricate relationships and dependencies that exist between users and items. To alleviate data scarcity effects, recent studies explored the potential of self-supervised learning to provide effective data augmentation. For example, in terms of contrastive augmentation, methods like SGL (Wu et al., 2021) and NCL (Lin et al., 2022) leverage graph contrastive learning to supplement the supervised recommendation task with the cross-view mutual information maximization. For generative augmentation, approaches such as AutoCF (Xia et al., 2023) are built upon the masked autoencoding mechanism, enabling the models to reconstruct the interaction structures for self-supervision.
While the recent advancements in self-supervised learning techniques have offered promising avenues for mitigating the impact of data scarcity in current collaborative filtering models, they also come with a significant limitation (Yuan et al., 2023). This limitation stems from the inherent design of the models, which heavily relies on using unique identities (IDs) to represent users and items throughout the entire process of learning representations. In practical recommenders, however, we often come across new recommendation data collected from different domains or time periods, involving diverse sets of users and items. This creates a challenge as existing ID-based recommendation models struggle to effectively incorporate and adapt to such new data, especially when there is a change in the set of identity tokens of users and items for zero-shot recommendation. The rigid dependence on user and item IDs in these models hinders their ability to generalize and perform well in scenarios where the user and item spaces are not static or fully overlapping.
For instance, the constant generation of new items like videos and social media content necessitates accurate recommendations in real-life recommender systems, even when there are limited interaction observations. While cross-domain recommendation methods (Xie et al., 2022; Cao et al., 2022) draw inspiration from leveraging information and knowledge across multiple domains to enhance recommendation performance, they often assume that users from different domains belong to the same set (Dacrema et al., 2012; Xie et al., 2022). Unfortunately, this assumption significantly restricts the flexibility and generalization capabilities of recommenders. As a result, existing methods may encounter difficulties in adapting and delivering accurate recommendations when confronted with diverse user populations across different domains.
@@
Language Models as Zero-Shot Recommenders. Considering the challenges and motivations discussed earlier, the objective of this study is to introduce a recommender system that functions as a zero-shot learner, possessing robust generalization capabilities and the flexibility to adapt to new recommendation data seamlessly. To accomplish this objective, we propose integrating language models with collaborative relation modeling, forming an effective text embedder-EasyRec that is both lightweight and effective. This integration seamlessly combines text-based semantic encoding with high-order collaborative signals, resulting in a recommender system that offers a strong model generalization ability.
@@
Recent research has explored leveraging large language models (LLMs) to enhance recommender systems. Existing approaches broadly fall into two categories. The first uses LLMs for data augmentation (e.g., RLMRec (Ren et al., 2024a), AlterRec (Li et al., 2024)), encoding textual information to complement collaborative filtering. While this combines LLM and collaborative strengths, these methods remain ID-based and struggle to generalize. The second approach utilizes LLMs to directly generate user-item interaction predictions based on language (e.g., LLaRA (Liao et al., 2024), CoLLM (Zhang et al., 2023)). However, such LLM-based recommenders face significant real-world limitations, suffering from poor efficiency (e.g., $\sim$1 second per prediction), rendering them impractical for large-scale recommendation tasks. These challenges highlight the need for more efficient, scalable solutions that seamlessly integrate the semantic understanding of textual information with the collaborative strengths of zero-shot recommenders.
@@
Our proposed model has demonstrated superior performance in comparison to state-of-the-art language models, as illustrated in Figure 1. This performance advantage is achieved within a cost-efficient parameter space, ranging from 100 to 400 million parameters, with the computational cost of $\sim$0.01 second per prediction. Notably, our model exhibits the scaling law phenomenon, where its performance continually improves as the parameter size is increased. Furthermore, in contrast to existing approaches that suffer from poor efficiency, our model is designed to be highly scalable and practical for large-scale recommendation tasks.
@@
Figure 1. EasyRecdemonstrates superior performance compared to state-of-the-art language models in text-based zero-shot recommendation on the Amazon and Steam datasets.
@@
Figure 2. The overall framework of our proposed collaborative information-guided language model EasyRec.
In summary, this work makes the following key contributions:
@@
  • Motivation. The primary objective of this study is to introduce a novel recommender system that is built upon language models to function as a zero-shot learner. This innovative approach aims to demonstrate exceptional adaptability to diverse recommendation data, while also exhibiting robust generalization capabilities.
@@
  • Methodology. To align text-based semantic encoding with collaborative signals from user behavior, we propose a novel contrastive learning-powered collaborative language modeling approach, which allows the system to capture both the semantic representations of users/items, as well as the underlying behavioral patterns and interactions within the recommendation data.
@@
  • Zero-Shot Recommendation Capacity. The EasyRec is extensively evaluated through rigorous experiments as a text-based zero-shot recommender. Performance comparisons reveal that it consistently achieves significant advantages over baseline methods in terms of recommendation accuracy and generalization capabilities. Furthermore, the study demonstrates the model’s remarkable potential in generating dynamic user profiles that are highly adaptive to time-evolving user preferences.
@@
  • Existing CF Model Enhancement. Our proposed EasyRec has been seamlessly integrated as a lightweight, plug-and-play component with state-of-the-art collaborative filtering (CF) models. The lightweight and modular design of our language model is a key strength, as it facilitates the adoption of our novel recommendation paradigm across a wide range of use cases.
2. Preliminaries
@@
In recommender systems, we have a set of users $\mathcal{U}$ and a set of items $\mathcal{I}$, along with the interactions between them (e.g., clicks, purchases). For each user $u\in\mathcal{U}$, we define $\mathcal{N}_{u}$ as the set of items that user $u$ has interacted with. Likewise, for each item $i\in\mathcal{I}$, we define $\mathcal{N}_{i}$ as the set of users who have interacted with that item. To represent these user-item interactions, we can use an interaction matrix $\mathcal{A}^{|\mathcal{U}|\times|\mathcal{I}|}$, where the entry $\mathcal{A}^{u,i}$ is 1 if user $u$ has interacted with item $i$, and 0 otherwise. The primary goal of a recommender model is to estimate the probability $p_{u,i}$ of a future interaction between a user $u$ and an item $i$. This predicted probability can then be used to generate personalized item recommendations for each user, tailored to their individual preferences and past behavior.
@@
Text-based Zero-Shot Recommendation is essential in recommender systems, as it can address the common cold-start problem. In real-world scenarios, new users and new items often lack sufficient interaction data, making it challenging to provide effective personalized recommendations. By leveraging the textual descriptions of users and items, such as product titles, details, and user profiles, the language models can construct semantic representations to enable text-based recommendations for these cold-start situations. This overcomes the limitations of traditional collaborative filtering methods, by offering a distinct advantage over the traditional ID-based paradigm. By leveraging robust language models, this approach exhibits remarkable potential in “zero-shot” scenarios, where the testing data have not been previously encountered.
@@
Formally, we define $\mathcal{P}_{u}$ and $\mathcal{P}_{i}$ as the generated text-based profiles of user $u$ and item $i$, respectively, which are encoded into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ using a language model, as shown as:
(1) $\mathbf{e}_{u}=\text{LM}(\mathcal{P}_{u}),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{e}_{i}=\text{LM}(\mathcal{P}_{i}).$
@@
The interaction preference score $p_{u,i}$ between a user $u$ and an item $i$ can be calculated as the cosine similarity between their text embeddings $\mathbf{e}_{u}$ and $\mathbf{e}i$, i.e., $p_{u,i}=cos(\mathbf{e}_{u},\mathbf{e}_{i})$. We can then recommend to the user the top-$k$ uninteracted items with the highest similarity scores, yielding a personalized recommendation set.
(2) $\mathcal{R}_{u}=\text{top\text{-}k}_{i\in\mathcal{I}\setminus\mathcal{N}_{u}}cos(\mathbf{e}_{u},\mathbf{e}_{i}).$
@@
Text-enhanced Collaborative Filtering. Collaborative filtering (CF) is a widely used recommender system approach that leverages the collaborative relationships among users and items. This existing CF paradigm can be enhanced by integrating encoded semantic representations. Typically, the preference score $p_{u,i}$ is calculated based on the interaction data, i.e., $p_{u,i}=f(u,i,\mathcal{A})$, where $\mathcal{A}$ represents the interaction data. Text-enhanced collaborative filtering builds upon this foundation by incorporating textual features $\mathbf{e}$ encoded by language models as supplementary representations. This integration aims to further improve the recommendation performance of the traditional ID-based collaborative filtering frameworks.
(3) $p_{u,i}=f(u,i,\mathcal{A},\mathbf{e}_{u},\mathbf{e}_{i}).$
3. Methodology
@@
In this section, we first discuss how we gather textual profiles for users and items in recommender systems, which are essential for pre-training and evaluating our model. Next, we’ll dive into the specifics of EasyRec and its training approach. Lastly, we’ll introduce our method for diversifying user profiles, which improves the model’s ability to adapt to various situations.
3.1. Collaborative User and Item Profiling
@@
In real-world recommenders, the only available information may be raw text data, such as product titles and categories, associated with the items. Privacy concerns often make it difficult to collect comprehensive user-side information. Furthermore, directly leveraging such textual information may overlook the crucial collaborative relationships needed for accurate user behavior modeling and preference understanding. To address these limitations, we propose to generate textual profiles by leveraging large language models (e.g., GPT, LLaMA series) (Ren et al., 2024a) to inject collaborative information into the textual profiles. This allows us to capture both the semantic and collaborative aspects of the items in a unified textual profile.
3.1.1. Item Profiling
@@
Given the raw item information, such as title $h_{i}$, categories $c_{i}$, and description $d_{i}$ (e.g., book summary), we aim to generate a comprehensive item profile $\mathcal{P}_{i}$ that captures both the semantic and collaborative aspects. To reflect user-item interactions, we incorporate the textual information (e.g., posted reviews $r_{u,i}$) provided by the item’s corresponding users. Formally, the item profile generation process is:
(4) $\mathcal{P}_{i}=\text{LLM}(\mathcal{M}_{i},h_{i},c_{i},\{r_{u,i}\})\,\,\text{or}\,\,\text{LLM}(\mathcal{M}_{i},h_{i},d_{i}),$
Here, $\mathcal{M}_{i}$ represents the generation instruction, while $h_{i}$ and $d_{i}$ serve as input. By leveraging large language models (LLMs), we can generate a concise yet informative item profile $\mathcal{P}_{i}$.
3.1.2. User Profiling

### v4

@@ Abstract
Deep neural networks have emerged as a powerful technique for learning representations from user-item interaction data in collaborative filtering (CF) for recommender systems. However, many existing methods heavily rely on unique user and item IDs, which restricts their performance in zero-shot learning scenarios. Inspired by the success of language models (LMs) and their robust generalization capabilities, we pose the question: How can we leverage language models to enhance recommender systems? We propose EasyRec, an effective approach that integrates text-based semantic understanding with collaborative signals. EasyRec employs a text-behavior alignment framework that combines contrastive learning with collaborative language model tuning. This ensures strong alignment between text-enhanced semantic representations and collaborative behavior information. Extensive evaluations across diverse datasets show EasyRec significantly outperforms state-of-the-art models, particularly in text-based zero-shot recommendation. EasyRec functions as a plug-and-play component that integrates seamlessly into collaborative filtering frameworks. This empowers existing systems with improved performance and adaptability to user preferences. Implementation codes are publicly available at: https://github.com/HKUDS/EasyRec.
EasyRec: Simple yet Effective Language Models for Recommendation
Xubin Ren The University of Hong Kong xubinrencs@gmail.com          Chao Huangthanks: Chao Huang is the Corresponding Author. The University of Hong Kong chaohuang75@gmail.com
1 Introduction
@@
Deep learning has established itself as a highly promising solution for capturing user preferences in online recommender systems Zhang et al. (2019); Yang et al. (2022); Zhang et al. (2022). This approach harnesses deep neural networks to learn rich user and item representations by analyzing complex user-item interaction patterns. Consequently, recommendation algorithms can accurately infer user preferences and deliver personalized recommendations Xu et al. (2023); Sun et al. (2021).
Recent advancements in enhancing recommender systems through neural network-powered collaborative filtering frameworks, particularly graph neural networks (GNNs) Wang et al. (2019); He et al. (2020); Xia et al. (2022), have leveraged the inherent graph structure in data to capture high-order relationships among users and items. Notable GNN-based approaches, such as NGCF Wang et al. (2019) and LightGCN He et al. (2020), demonstrate impressive performance via recursive message passing mechanisms. However, data scarcity remains a significant challenge, hindering deep collaborative filtering models from accurately learning user/item representations, especially with sparse interaction data Lin et al. (2021); Wei et al. (2021); Hao et al. (2021). To address this, recent studies have explored self-supervised learning for effective data augmentation. For instance, contrastive methods like SGL Wu et al. (2021) and NCL Lin et al. (2022) utilize graph contrastive learning, while generative approaches like AutoCF Xia et al. (2023) employ masked autoencoding to reconstruct interaction structures.
Recent advances in self-supervised learning show promise in alleviating data scarcity in collaborative filtering models. However, these methods encounter a significant limitation Yuan et al. (2023). They rely heavily on unique identities (IDs) to represent users and items. This reliance presents major challenges for practical recommenders that must handle data from diverse domains or time periods effectively. Existing ID-based models struggle to adapt to changes in user and item identity tokens. This problem becomes particularly severe in zero-shot recommendation scenarios Ding et al. (2021); Hou et al. (2022). In these cases, training data lacks overlap with deployment data regarding users and items. This limitation hinders their ability to generalize effectively towards foundational recommender systems. Cross-domain recommendation methods Xie et al. (2022); Cao et al. (2022) attempt to leverage knowledge across multiple domains. However, they often make restrictive assumptions about user overlap. These approaches assume that users from different domains belong to the same set Dacrema et al. (2012); Xie et al. (2022). This assumption significantly limits their flexibility and generalization capabilities. Consequently, existing methods struggle to provide accurate recommendations to diverse user populations. They face particular challenges when operating across varying domains and contexts.
@@
Language Models as Zero-Shot Recommenders. The challenges discussed earlier highlight a critical need in recommendation systems. This study aims to introduce a recommender system that functions as a zero-shot learner. The system should possess robust generalization capabilities and adapt to new recommendation data seamlessly. To accomplish this objective, we propose integrating language models with collaborative relation modeling. This integration forms an effective text embedder-EasyRec that is both lightweight and effective. Our approach seamlessly combines text-based semantic encoding with high-order collaborative signals. This combination results in a recommender system with strong generalization ability. The system leverages rich semantic understanding while capturing collaborative patterns from user-item interactions.
@@
Recent research has explored leveraging large language models (LLMs) to enhance recommender systems. Existing approaches broadly fall into two categories. The first category uses LLMs for data augmentation (e.g., RLMRec Ren et al. (2024), AlterRec Li et al. (2024)). These methods encode textual information to complement collaborative filtering. While this combines LLM and collaborative strengths, these methods remain ID-based and struggle to generalize. The second approach utilizes LLMs to directly generate user-item interaction predictions (e.g., LLaRA Liao et al. (2024), CoLLM Zhang et al. (2023)). However, such LLM-based recommenders suffer from poor efficiency, requiring approximately one second per prediction. This renders them impractical for large-scale recommendation tasks. These challenges highlight the need for efficient, scalable solutions that integrate semantic understanding with collaborative strengths for zero-shot recommendation.
@@
Our model demonstrates superior performance compared to state-of-the-art language models, as illustrated in Figure 1. This performance advantage is achieved within a cost-efficient parameter space of 100 to 400 million parameters. The computational cost is approximately 0.01 seconds per prediction. Notably, our model exhibits the scaling law phenomenon. Performance continually improves as parameter size increases. This contrasts with existing approaches that suffer from poor efficiency. Our model is designed to be highly scalable and practical for large-scale recommendation tasks. The computational efficiency represents significant advancement over current methods. While existing LLM-based recommenders require substantial inference time, our approach maintains both accuracy and speed for real-world deployment. In summary, this work makes the following contributions:
@@
Figure 1: EasyRec outperforms state-of-the-art language models in text-based zero-shot recommendation.
@@
Figure 2: The overall framework of our proposed collaborative information-guided language model EasyRec.
@@
  • Motivation. The primary objective of this study is to introduce a novel recommender system built upon language models. This system functions as a zero-shot learner with exceptional adaptability to diverse recommendation data. The new EasyRec exhibits robust generalization capabilities across different domains and contexts.
@@
  • Methodology. We propose a novel contrastive learning-powered collaborative language modeling approach. This method aligns text-based semantic encoding with collaborative signals from user behavior. The system captures both semantic representations of users and items. It also learns underlying behavioral patterns and interactions within the recommendation data.
@@
  • Zero-Shot Recommendation Capacity. The EasyRec is extensively evaluated through rigorous experiments as a text-based zero-shot recommender system. Performance comparisons reveal consistent and significant advantages over baseline methods. The model excels in both recommendation accuracy and generalization capabilities. Furthermore, the study demonstrates remarkable potential in adapting dynamic user profiles. These profiles are highly adaptive to time-evolving user preferences.
@@
  • Existing Recommender Enhancement. Our proposed EasyRec integrates seamlessly as a lightweight, plug-and-play component with state-of-the-art collaborative filtering models. The lightweight and modular design represents a key strength of our approach. This design facilitates adoption of our novel recommendation paradigm across diverse use cases. The integration enhances existing recommendation methods without significant computational overhead.
2 Preliminaries
@@
In recommender systems, we have a set of users $\mathcal{U}$ and items $\mathcal{I}$, along with their interactions (e.g., clicks, purchases). For each user $u\in\mathcal{U}$, let $\mathcal{N}_{u}$ be the set of items interacted with. For each item $i\in\mathcal{I}$, let $\mathcal{N}_{i}$ be the set of users who have interacted with it. These interactions can be represented by an interaction matrix $\mathcal{A}^{|\mathcal{U}|\times|\mathcal{I}|}$. Here, $\mathcal{A}^{u,i}$ is 1 if user $u$ has interacted with item $i$, and 0 otherwise. The goal of recommendation is to predict the preference score $p_{u,i}$ of future interactions between user $u$ and item $i$. This score can be used to generate recommendations based on individual preferences.
@@
Text-based Zero-Shot Recommendation is crucial in recommender systems. It serves as a pathway toward foundation models. Textual descriptions, such as product titles and user profiles, provide valuable information. These descriptions enable effective recommendations across various datasets. This approach overcomes the limitations of traditional collaborative filtering methods. It offers significant advantage in generalization over the ID-based paradigm. This capability extends to scenarios where specific user-item interactions have not been previously encountered.
@@
Formally, we define $\mathcal{P}_{u}$ and $\mathcal{P}_{i}$ as the generated text-based profiles of user $u$ and item $i$, respectively. These profiles are then encoded into representations $\mathbf{e}_{u}$ and $\mathbf{e}_{i}$ using a language model (LM). This process is shown as:
$\mathbf{e}_{u}=\text{LM}(\mathcal{P}_{u}),~~~\mathbf{e}_{i}=\text{LM}(\mathcal{P}_{i}).$ (1)
@@
The preference score $p_{u,i}$ between user $u$ and item $i$ is calculated as the cosine similarity between their text embeddings $\mathbf{e}_{u}$ and $\mathbf{e}i$. This is expressed as $p_{u,i}=\cos(\mathbf{e}_{u},\mathbf{e}_{i})$. We then recommend the top-$k$ unclicked items with highest similarity scores, resulting in a recommendation set.
$\mathcal{R}_{u}=\text{top\text{-}k}_{i\in\mathcal{I}\setminus\mathcal{N}_{u}}cos(\mathbf{e}_{u},\mathbf{e}_{i}).$ (2)
@@
Text-enhanced Collaborative Filtering. Collaborative filtering (CF) is a widely used recommendation paradigm. It leverages the collaborative relationships among users and items. This existing CF paradigm can be enhanced by integrating encoded semantic representations. Typically, the value $p_{u,i}$ is calculated based on the interaction data. This is expressed as $p_{u,i}=f(u,i,\mathcal{A})$, where $\mathcal{A}$ represents the interaction data. Text-enhanced collaborative filtering builds upon this foundation. It incorporates textual features $\mathbf{e}$ encoded by language models as supplementary representations. This integration aims to improve the recommendation performance of traditional ID-based frameworks.
$p_{u,i}=f(u,i,\mathcal{A},\mathbf{e}_{u},\mathbf{e}_{i}).$ (3)
3 Methodology
@@
In this section, we first discuss how we gather textual profiles for users and items. Next, we dive into the specifics of EasyRec and its training approach. Lastly, we introduce our method for diversifying user profiles. This method improves the model’s ability to adapt to various situations.
3.1 Collaborative User and Item Profiling
@@
In real-world recommenders, only raw text data may be available. This includes item titles and categories. Privacy concerns limit comprehensive user information. Directly using this textual data can overlook essential collaborative relationships. These relationships are needed for effective user behavior modeling. To address these issues, we propose generating textual profiles using large language models. Examples include GPT and LLaMA series Ren et al. (2024). These models incorporate collaborative information.
3.1.1 Item Profiling
@@
Given raw item information, we aim to generate a comprehensive item profile $\mathcal{P}_{i}$. This information includes title $h_{i}$, categories $c_{i}$, and description $d_{i}$ (e.g., book summary). The profile captures both semantic and collaborative aspects. To reflect user-item interactions, we incorporate textual information. This includes user reviews $r_{u,i}$. The item profile generation process is:
$\mathcal{P}_{i}=\text{LLM}(\mathcal{M}_{i},h_{i},c_{i},\{r_{u,i}\})\vee\text{LLM}(\mathcal{M}_{i},h_{i},d_{i}),$ (4)
Here, $\mathcal{M}_{i}$ depicts the generation instruction. We consider two scenarios. One includes a description in the raw data (right-hand side). The other does not include a description (left-hand side). In the latter case, we utilize collaborative reviews to inform the profile. Using LLMs, we can generate an informative item profile $\mathcal{P}_{i}$.
3.1.2 User Profiling